OpenStudy (anonymous):

Trigonometry Establish the identity cot^2 u/2 = csc u + cot u/csc u - cot u

OpenStudy (sidsiddhartha):

\[\tan u/2 =\frac{ sinu }{ 1+cosu }\] \[\tan ^{2}u/2=\frac{ \sin ^{2} }{ (1+\cos ^{}u)^{2} }=\frac{ \sin ^{2}u }{ (1+2cosu+\cos ^{2}u )}\]

OpenStudy (sidsiddhartha):

now \[\cot ^{2}u=\frac{ 1 }{ \tan ^{2}u }\] use this